OpenStudy (anonymous):

Find the length of the curve defined by y = 6x^(3/2) + 7 from x = 4 to x = 7.

OpenStudy (anonymous):

haha, Same goes with that question.

OpenStudy (anonymous):

help bro

OpenStudy (anonymous):

first derivative the function with respect to y. \[\frac{ 18 }{ 2 }\times x ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

simplify, 9 sqrt(x)

OpenStudy (anonymous):

why is it x^1/2

OpenStudy (anonymous):

oh cause -1?

OpenStudy (anonymous):

to derivative, exponential functions, bring down the original exponent, then minus one the exponent to find the new exponent

OpenStudy (anonymous):

\[\int_4^7\sqrt{1+(f'(x))^2}dx\]

OpenStudy (anonymous):

i need the steps though

OpenStudy (anonymous):

the problem worked out

OpenStudy (anonymous):

\[f(x)=6x^{\frac{3}{2}}\] \[f'(x)=9x^{\frac{1}{2}}\] \[(f'(x))^2=81x\] \[\int_4^7\sqrt{1+81x}dx\]

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

integrate?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i need help though

OpenStudy (anonymous):

try \(u=1+81x\)

OpenStudy (anonymous):

whats that suppse to do

OpenStudy (anonymous):

\[\int\limits_{4}^{7}\sqrt{1+81x}dx\]

OpenStudy (anonymous):

but what do i do from there

OpenStudy (anonymous):

let 1+81x = u then du = 81dx

OpenStudy (anonymous):

substitute

OpenStudy (anonymous):

ok so...

OpenStudy (anonymous):

\[\frac{ 1 }{ 81 } \int\limits_{4}^{7}u ^{\frac{ 1 }{ 2 }}du\]

OpenStudy (anonymous):

why did you pull out 1/81 and was it u^3/2+1

OpenStudy (anonymous):

? help

OpenStudy (anonymous):

look again from the original equation, you should substitute dx with du/81 right?

OpenStudy (anonymous):

can u draw it

OpenStudy (anonymous):

ist't it that you have the integral of the sqrt of (1 + 81x) dx? look above

OpenStudy (anonymous):

yes the part that satelitte drew

OpenStudy (anonymous):

thats where im at, can u work with me from there

OpenStudy (anonymous):

ok, then you let 1+81x = u got it?

OpenStudy (anonymous):

then du is going to be equal to 81dx got it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok from there, dx = du/81 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then substitute you will arrive with my final equation a while ago got it?

OpenStudy (anonymous):

can u show me

OpenStudy (anonymous):

?

OpenStudy (anonymous):

replace 1+81x with u then dx with du/81

OpenStudy (anonymous):

so its sqrt(1+81x) * 81 dx?

OpenStudy (anonymous):

its going to be sqrt(u) du / 81

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

then you can now integrate.

OpenStudy (anonymous):

i thought we just did , can u just stay on my problem and help me

OpenStudy (anonymous):

so you're going to have this final equation after substitution \[\frac{ 1 }{ 81 } \int\limits_{4}^{7}\sqrt{u}\]

OpenStudy (anonymous):

du

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

integrate you will arrive at \[u ^{\frac{ 1 }{ 2 }+1} \div (\frac{ 1 }{2 }+1) \] evaluated from 4 to 7

OpenStudy (anonymous):

where did the 1/81 go?

OpenStudy (anonymous):

nevermind so now what

OpenStudy (anonymous):

i just forgot to type it, but it will stay there

OpenStudy (anonymous):

ok i got it 63?

OpenStudy (anonymous):

\[\frac{ 1 }{ 81 } u ^{\frac{ 3 }{ 2 }}\times \frac{ 1 }{ \frac{ 3 }{ 2 } }\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

there you go.

OpenStudy (anonymous):

63